# Preconditioned log-determinant approximation:

### v1

@@ 5.5 Comparison with existing strategies
For a fixed budget of matvecs, we compare our one-sample strategy (one-sample) and log-det-ective strategy (Algorithm 1) with several existing techniques:
@@
  • the strategy ($\alpha$-rank) which splits the number of matvecs evenly between the creation of the preconditioner. This is called “half samples” in the plots and is inspired by [WengerPleissHennigCunninghamGardner22, PerssonKressner23].
Motivated by the results in Section 5.3, we fix the number of iterations of the Lanczos method as $m=10$ in all the methods where a preconditioner is used. When comparing with (SLQ), however, we should run the Lanczos algorithm for a potentially larger number of steps $\widehat{m}$; we choose the smallest $\widehat{m}$ such that
@@
As before, for each value of $\ell$ that we consider, we run each algorithm $100$ times; the plots in Figure 7 report the mean relative error, with respect to the target quantity ${\rm trace}\log(A+I)$, and the standard deviation obtained discarding the best and the worst $10\%$ errors over the runs. The $x$-axis is the total number of matvecs, the $y$-axis is the final relative error in the computation of the log-determinant and the vertical bars denote the standard deviation. Figure 7 confirms that the one-sample strategy (one-sample) is almost always close to optimal. Moreover, the log-det-ective method (Algorithm 1) is very effective, and it is able to correctly identify the cases in which the one-sample strategy is suboptimal.
Figure 7: Comparison of the errors of the strategies listed in Section 5.5, considering a Nyström approximation of dimension $\ell=100,\ldots,1000$ and $m=10$ steps of the Lanczos method for the methods involving a preconditioner and number of steps $\widehat{m}$ that satisfies (34) for the strategy (SLQ), averaging on $100$ runs.
@@ 6 Conclusions
We proposed two variance reduction methods to approximate the regularized log-determinant of an SPSD matrix $A$, which combine preconditioning and the use of few samples for estimating the residual term via SLQ. We used a generalized Nyström preconditioner, which is cheap to compute and effective for matrices with some spectral decay. We theoretically analyzed the method in the case when we pick one single sample for SLQ, and we tested this strategy on matrices exhibiting a variety of eigenvalue decays, showing numerically that this choice is often the optimal one. We then proposed a log-det-ective strategy (Algorithm 1) which detects when the convergence of the one-sample strategy is not close to optimal and, in that case, changes the allocation of the budget of matvecs, allowing more vectors for the SLQ part. Several numerical experiments justify the choice of parameters for the log-det-ective method and confirmed the effectiveness of both strategies.

### v2

@@ 5.5 Comparison with existing strategies
For a fixed budget of matvecs, we compare our one-sample strategy (one-sample) and log-det-ective strategy (Algorithm 1) with several existing techniques that are tailored to the computation of traces of matrix functions:
@@
  • the funNys++ algorithm in [PerssonKressner23];
  • the Krylov-aware method, corresponding to [ChenHallman23, Algorithm 3.1];
@@
  • the strategy ($\alpha$-rank) which splits the number of matvecs evenly between the creation of the preconditioner. This is called “half samples” in the plots and is inspired by [WengerPleissHennigCunninghamGardner22].
Motivated by the results in Section 5.3, we fix the number of iterations of the Lanczos method as $m=10$ in all the methods where a preconditioner (either in an implicit or explicit way) is used. In particular, for the Krylov-aware strategy [ChenHallman23] we need to set a block size $b$, a deflation parameter $q$ which represents the number of blocks used for the implicit preconditioner, and a number of samples $N$ to estimate the (implicit) residual. In our experiment, we set $b=4$, $N=1$, and $q=\max\left\{0,\left\lfloor\frac{(\ell+m)-(b+N)m}{b}\right\rfloor\right\}$, so that we use at least $\ell+m$ matvecs, which is our budget. The rationale behind this choice is the same used for our proposed strategies: we are (implicitly) preconditioning as much as we can and use few matvecs to estimate the (implicit) residual. In this sense, the Krylov-aware technique can be seen as a different preconditioning strategy. When running the algorithm, in some examples we stop before the budget reaches $1010$ matvecs because the deflated space becomes numerically rank-deficient.
When comparing with (SLQ) and funNys++, however, we should run the Lanczos algorithm for a potentially larger number of steps $\widehat{m}$; we choose the smallest $\widehat{m}$ such that
@@
As before, for each value of $\ell$ that we consider, we run each algorithm $100$ times; the plots in Figure 7 report the mean relative error, with respect to the target quantity ${\rm trace}\log(A+I)$, and the standard deviation obtained discarding the best and the worst $10\%$ errors over the runs. The $x$-axis is the total number of matvecs, the $y$-axis is the final relative error in the computation of the log-determinant and the vertical bars denote the standard deviation. Again for clarity purposes, the lines related to the proposed strategies are thicker, and others are slightly shifted to the left and to the right, although the budget of matvecs used for different methods is still the same. Figure 7 confirms that the one-sample strategy (one-sample) is almost always close to optimal, and generally better than splitting the budget equally (the “half-samples” strategy). As expected, SLQ is slow to converge since it does not exploit any low-rank structure in the matrix $A$; on the other side, the strategy (low-rank) does not take into account the residual an also performs rather suboptimally in most examples. As the Krylov-aware algorithm is also building an implicit preconditioner and we chose the parameters specifically to mimic a large preconditioner and a small budget for the implicit residual, its performance is similar to our proposed methods for larger values of the budget, while it sometimes has a slower start.
Finally, the log-det-ective method (Algorithm 1) is very effective, and it is able to correctly identify the cases in which the one-sample strategy is suboptimal.
Figure 7: Comparison of the errors of the strategies listed in Section 5.5, considering a Nyström approximation of dimension $\ell=100,\ldots,1000$ and $m=10$ steps of the Lanczos method for the methods involving a preconditioner and number of steps $\widehat{m}$ that satisfies (34) for the strategy (SLQ) and funNys++, averaging on $100$ runs.
@@ 6 Conclusions
We proposed two variance reduction methods to approximate the regularized log-determinant of an SPSD matrix $A$, which combine preconditioning and the use of few samples for estimating the residual term via SLQ. We used a generalized Nyström preconditioner, which is cheap to compute and effective for matrices with some spectral decay. We theoretically analyzed the method in the case when we pick one single sample for SLQ, and we tested this strategy on matrices exhibiting a variety of eigenvalue decays, showing numerically that this choice is often the optimal one. For matrices with moderate spectral decay, we observe that, although suboptimal, it still comparable to the best methods, and thus it is a safe simple choice. To automatically detect the suboptimal situations, we proposed a “log-det-ective” strategy (Algorithm 1) which, in these cases, changes the allocation of the budget of matvecs, allowing more vectors for the SLQ part. Several numerical experiments justify the choice of parameters for the “log-det-ective” method and confirmed the effectiveness of both strategies.